OpenStudy (anonymous):

Switching to which base would allow you to easily reduce this down to the fraction 3/2 ? log^4(8)

OpenStudy (anonymous):

is this your question Switching to which base would allow you to easily reduce this down to the fraction 3/2 ? log_4(8) a) base 8 b) base 4 c) base 2 d) base 10

OpenStudy (anonymous):

cause if so the answer would be c

OpenStudy (anonymous):

since 4 = 2^2 and 8 = 2^3

OpenStudy (anonymous):

yes thankyou ! how did you do it like solve it ?

OpenStudy (anonymous):

It was a problem i asked before so i just took the answer that they gave me.

OpenStudy (anonymous):

They didn't really say how to solve it.

OpenStudy (jdoe0001):

soooooo

OpenStudy (anonymous):

Well hoped this helped.

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

oh lol thanks tho (: it did !

OpenStudy (jdoe0001):

aheem... I'd think not how is guessing help?

OpenStudy (anonymous):

buh byee

OpenStudy (anonymous):

it doesnt but i am just going to ask my online tutor how to explain it to me (:

OpenStudy (ybarrap):

The only way you can get 3/2 from a base of 2 is if $$ \cfrac{\log_?8}{\log_?4}=\cfrac{\log_22^3}{\log_22^2}=\cfrac{3\log_22}{2\log_22}=\cfrac{3}{2} $$